# Minimal projections onto spaces of polynomials on real euclidean spheres

### v3

@@ 1. Introduction
It is worth noting that projections play an important role in approximation theory. In fact, if $Y$ is a Banach space and $P$ a projection from $Y$ onto a subspace $X$, then the error $\|y-Py\|_{Y}$ of approximation of an element $y\in Y$ by $Py$ satisfies
@@
The initial inspiration for this work stems from the seminal contributions of Ryll and Wojtaszczyk [34]. Solving a problem posed by Waigner, they demonstrated that the inclusion $H_{\infty}(B_{\ell_{2}^{n}(\mathbb{C})})\hookrightarrow H_{1}(B_{\ell_{2}^{n}(\mathbb{C})})$ between Hardy spaces is not compact, where $B_{\ell_{2}^{n}(\mathbb{C})}$ denotes the open unit ball of the complex $n$-dimensional Hilbert space $\ell_{2}^{n}(\mathbb{C})$ and $n>1$. As noted in [34], this result is intimately related to a question raised by Rudin in his monograph [32]: Does there exist an inner function on the open unit ball of the Hilbert space $\ell_{2}^{n}(\mathbb{C})$, $n>1$ ? Recall that a nonconstant bounded holomorphic function with domain $B_{\ell_{2}^{n}(\mathbb{C})}$ is called an inner function if its radial limits have modulus equal to $1$ almost everywhere on $\mathbb{S}^{n-1}$, where ”almost everywhere” refers to the rotation-invariant probability measure on $\mathbb{S}^{n-1}$. For more details, we refer the reader to [31] and [37].
@@ 2. Preliminaries
We denote the Lebesgue measure on $\mathbb{R}^{n}:=\ell_{2}^{n}(\mathbb{R})$ by $\lambda_{n}$, and the surface (Borel) measure on the real unit sphere $\mathbb{S}^{n-1}$ in $\ell_{2}^{n}(\mathbb{R})$ by $s_{n}$. Recall one of the simplest definitions of $s_{n}$: If $A$ is a Borel set in $\mathbb{S}^{n-1}$, then
@@ 3.1. Averaging projections
Clearly, $\phi_{id}$ is the identity on $C(\mathbb{S}^{n-1})$ (where $id$ denotes the unity in $\mathcal{O}_{n}$), and $\phi_{AB}=\phi_{B}\phi_{A}$ for all $A,B\in\mathcal{O}_{n}$. Furthermore,
@@ 3.2. Axially invariance
First, we need to establish some notation. Let $\mathcal{O}_{n}(e_{1})$ denote the subgroup of orthogonal transformations that fix $e_{1}$; that is, $A\in\mathcal{O}_{n}(e_{1})$ if $A$ is orthogonal and $Ae_{1}=e_{1}$. We consider functions $f:\mathbb{R}^{n}\to\mathbb{C}$ that are invariant under the action of this subgroup, meaning that $f\circ A=f$ for every $A\in\mathcal{O}_{n}(e_{1})$. We will refer to such functions, somewhat informally, as axially invariant.
@@
We write $x=(x_{1},x^{\prime})\in\mathbb{R}\times\mathbb{R}^{n-1}$ for $x\in\mathbb{R}^{n}$, and note that every $n\times n$-matrix $A$ belongs to $\mathcal{O}_{n}(e_{1})$ if and only if it has the form
@@
3.3. Integral representations
As announced our main intention in this article is to collect information on the projection constants of $\mathcal{H}_{d}(\mathbb{S}^{n-1})$, $\mathcal{P}_{d}(\mathbb{S}^{n-1})$ and $\mathcal{P}_{\leq d}(\mathbb{S}^{n-1})$. The following theorem follows directly from Theorem 3.1 and Proposition 3.3.
@@ 4.1. Harmonics
We start with the calculation of the projection constant of $\mathcal{H}_{d}(\mathbb{S}^{1})$, so the dimension $n$ equals $2$.
@@ 4.2. Homogeneous polynomials
As before in Proposition 4.2 we start with the $2$-dimensional case.
@@
It should be pointed out that in the case of the complex Hilbert space $\ell_{2}^{2}(\mathbb{C})$, the projection constant $\boldsymbol{\lambda}\big(\mathcal{P}_{d}(\ell_{2}^{2}(\mathbb{C})\big)$ is bounded by $2$ for any degree $d$ (see again (3) and also (4)). However, the previous result shows that for the real case we have a significant difference. In fact, the projection constant of $\mathcal{P}_{d}(\ell_{2}^{2}(\mathbb{R}))=\mathcal{P}_{d}(\mathbb{S}^{1})$ has a logarithmic increase with respect to $d$.
@@
$\lim_{d\to\infty}\frac{\boldsymbol{\lambda}\big(\mathcal{P}_{d}(\mathbb{S}_{n}^{\mathbb{R}})\big)}{d^{\frac{n-2}{2}}}$ $=\,\frac{1}{2\sqrt{\pi}}\frac{\Gamma(\frac{n}{2})}{\,\Gamma(n-1)}\,\left[\lim_{d\to\infty}\,\,\frac{\Gamma(d+n)}{\Gamma(d+\frac{n+1}{2})d^{\frac{n-1}{2}}}\right]\,\left[\lim_{d\to\infty}\,2\sqrt{d}\,\int_{0}^{1}\big|P_{d}^{(\frac{n-1}{2},\frac{n-1}{2})}(t)\big|(1-t^{2})^{\frac{n-3}{2}}dt\right]$
@@
$\lim_{d\to\infty}\frac{\boldsymbol{\lambda}\big(\mathcal{P}_{d}(\mathbb{S}_{n}^{\mathbb{R}})\big)}{d^{\frac{n-2}{2}}}$ $=\,\frac{2^{n-1}}{\pi^{2}\Gamma(n-1)\Gamma(\frac{n}{4}-\frac{1}{2})}\,\frac{\sqrt{\pi}\Gamma(\frac{n}{2})}{2^{\frac{n}{2}-1}\Gamma(\frac{n}{4})}\,\,$
@@ 4.3. Finite degree polynomials
Indeed, this easily follows by combining the Equations (15), (17) and (43). Hence, as in the case of Corollary 4.7 (see also Remark 4.8), this again shows that here the Kadets-Snobar theorem leads to a weak upper estimate - in contrast to the result for homogeneous spherical harmonics from Theorem 4.4.
@@
$\lim_{d\to\infty}\frac{\boldsymbol{\lambda}\big(\mathcal{P}_{\leq d}(\mathbb{S}_{n}^{\mathbb{R}})\big)}{d^{\frac{n-2}{2}}}$
@@ 5. Transition to Real Coefficients
Nevertheless, the Banach-space geometry of the underlying spaces may change. A striking example is provided by the degree-$d$ harmonics on the circle. As noted in the comments following Proposition 4.2, the complex version $\mathcal{H}_{d}(\mathbb{S}^{1})$ is isometrically isomorphic to the complex space $\ell_{1}^{2}(\mathbb{C})$, whereas its real-coefficient counterpart turns out to be a $2$-dimensional real Hilbert space. We give a short proof of this fact.
@@
is an isometric linear bijection of real Banach spaces. In particular, the real version of $\mathcal{H}_{d}(\mathbb{S}^{1})$ is a $2$-dimensional Hilbert space.

### v4

@@ 1. Introduction
It is worth noting that projections play an important role in approximation theory. In fact, if $Y$ is a Banach space and $P$ a projection from $Y$ onto a subspace $X$, then the error $\|y-Py\|_{Y}$ of approximation of an element $y\in Y$ by PyPy satisfies
@@
The initial inspiration for this work stems from the seminal contributions of Ryll and Wojtaszczyk [34]. Solving a problem posed by Waigner, they demonstrated that the inclusion $H_{\infty}(B_{\ell_{2}^{n}(\mathbb{C})})\hookrightarrow H_{1}(B_{\ell_{2}^{n}(\mathbb{C})})$ between Hardy spaces is not compact, where $B_{\ell_{2}^{n}(\mathbb{C})}$ denotes the open unit ball of the complex $n$-dimensional Hilbert space $\ell_{2}^{n}(\mathbb{C})$ and $n>1$. As noted in [34], this result is intimately related to a question raised by Rudin in his monograph [32]: Does there exist an inner function on the open unit ball of the Hilbert space $\ell_{2}^{n}(\mathbb{C})$, $n>1$ ? Recall that a nonconstant bounded holomorphic function with domain $B_{\ell_{2}^{n}(\mathbb{C})}$ is called an inner function if its radial limits have modulus equal to 11 almost everywhere on $\mathbb{S}^{n-1}$, where ”almost everywhere” refers to the rotation-invariant probability measure on $\mathbb{S}^{n-1}$. For more details, we refer the reader to [31] and [37].
@@ 2. Preliminaries
We denote the Lebesgue measure on $\mathbb{R}^{n}:=\ell_{2}^{n}(\mathbb{R})$ by $\lambda_{n}$, and the surface (Borel) measure on the real unit sphere $\mathbb{S}^{n-1}$ in $\ell_{2}^{n}(\mathbb{R})$ by $s_{n}$. Recall one of the simplest definitions of $s_{n}$: If AA is a Borel set in $\mathbb{S}^{n-1}$, then
@@ 3.1. Averaging projections
Clearly, $\phi_{id}$ is the identity on $C(\mathbb{S}^{n-1})$ (where idid denotes the unity in $\mathcal{O}_{n}$), and $\phi_{AB}=\phi_{B}\phi_{A}$ for all $A,B\in\mathcal{O}_{n}$. Furthermore,
@@ 3.2. Axially invariance
First, we need to establish some notation. Let $\mathcal{O}_{n}(e_{1})$ denote the subgroup of orthogonal transformations that fix $e_{1}$; that is, $A\in\mathcal{O}_{n}(e_{1})$ if AA is orthogonal and $Ae_{1}=e_{1}$. We consider functions $f:\mathbb{R}^{n}\to\mathbb{C}$ that are invariant under the action of this subgroup, meaning that $f\circ A=f$ for every $A\in\mathcal{O}_{n}(e_{1})$. We will refer to such functions, somewhat informally, as axially invariant.
@@
We write $x=(x_{1},x^{\prime})\in\mathbb{R}\times\mathbb{R}^{n-1}$ for $x\in\mathbb{R}^{n}$, and note that every $n\times n$-matrix AA belongs to $\mathcal{O}_{n}(e_{1})$ if and only if it has the form
@@
$3.$ Integral representations
As announced our main intention in this article is to collect information on the projection constants of $\mathcal{H}_{d}(\mathbb{S}^{n-1})$, $\mathcal{P}_{d}(\mathbb{S}^{n-1})$ and $\mathcal{P}_{\leq d}(\mathbb{S}^{n-1})$. The following theorem follows directly from Theorem 3.1 and Proposition $3.$
@@ 4.1. Harmonics
We start with the calculation of the projection constant of $\mathcal{H}_{d}(\mathbb{S}^{1})$, so the dimension $n$ equals 22.
@@ 4.2. Homogeneous polynomials
As before in Proposition 4.2 we start with the 22-dimensional case.
@@
It should be pointed out that in the case of the complex Hilbert space $\ell_{2}^{2}(\mathbb{C})$, the projection constant $\boldsymbol{\lambda}\big(\mathcal{P}_{d}(\ell_{2}^{2}(\mathbb{C})\big)$ is bounded by 22 for any degree $d$ (see again (3) and also (4)). However, the previous result shows that for the real case we have a significant difference. In fact, the projection constant of $\mathcal{P}_{d}(\ell_{2}^{2}(\mathbb{R}))=\mathcal{P}_{d}(\mathbb{S}^{1})$ has a logarithmic increase with respect to $d$.
@@
$\lim_{d\to\infty}\frac{\boldsymbol{\lambda}\big(\mathcal{P}_{d}(\mathbb{S}^{n-1})\big)}{d^{\frac{n-2}{2}}}$ $=\,\frac{1}{2\sqrt{\pi}}\frac{\Gamma(\frac{n}{2})}{\,\Gamma(n-1)}\,\left[\lim_{d\to\infty}\,\,\frac{\Gamma(d+n)}{\Gamma(d+\frac{n+1}{2})d^{\frac{n-1}{2}}}\right]\,\left[\lim_{d\to\infty}\,2\sqrt{d}\,\int_{0}^{1}\big|P_{d}^{(\frac{n-1}{2},\frac{n-1}{2})}(t)\big|(1-t^{2})^{\frac{n-3}{2}}dt\right]$
@@
$\lim_{d\to\infty}\frac{\boldsymbol{\lambda}\big(\mathcal{P}_{d}(\mathbb{S}^{n-1})\big)}{d^{\frac{n-2}{2}}}$ $=\,\frac{2^{n-1}}{\pi^{2}\Gamma(n-1)\Gamma(\frac{n}{4}-\frac{1}{2})}\,\frac{\sqrt{\pi}\Gamma(\frac{n}{2})}{2^{\frac{n}{2}-1}\Gamma(\frac{n}{4})}\,\,$
@@ 4.3. Finite degree polynomials
Indeed, this easily follows by combining the Equations (15), (17) and (43). Hence, as in the case of Corollary 4.7 (see also Remark 4.8), this again shows that here the Kadets-Snobar theorem leads to a weak upper estimate - in contrast to the result for homogeneous spherical harmonics from Theorem $4.$
@@
$\lim_{d\to\infty}\frac{\boldsymbol{\lambda}\big(\mathcal{P}_{\leq d}(\mathbb{S}^{n-1})\big)}{d^{\frac{n-2}{2}}}$
@@ 5. Transition to Real Coefficients
Nevertheless, the Banach-space geometry of the underlying spaces may change. A striking example is provided by the degree-$d$ harmonics on the circle. As noted in the comments following Proposition 4.2, the complex version $\mathcal{H}_{d}(\mathbb{S}^{1})$ is isometrically isomorphic to the complex space $\ell_{1}^{2}(\mathbb{C})$, whereas its real-coefficient counterpart turns out to be a 22-dimensional real Hilbert space. We give a short proof of this fact.
@@
is an isometric linear bijection of real Banach spaces. In particular, the real version of $\mathcal{H}_{d}(\mathbb{S}^{1})$ is a 22-dimensional Hilbert space.